y= x + 3
y= 4x -3

solve each system using substitution

Answer by stanbon(75887)

y= x + 3
y= 4x -3
----------
x+3 = 4x-3
3x=6
x = 2
-------
Substitute into y = x+3 to solve for y:
y= 2+3
y = 5
============
Cheers,
Stan H.